OpenStudy (anonymous):

K is the midpoint of line AB . if line AK= 3x + 1 and line KB is 1/2x+6 , find the length of AK

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

put it in equation form you know how to do that right?

OpenStudy (anonymous):

nope

OpenStudy (lgbasallote):

lol AK has already been given :P copy it

OpenStudy (anonymous):

since k bisects the line

OpenStudy (anonymous):

dont i have to solve forit

OpenStudy (anonymous):

that big line is is k (forgot to label it)

OpenStudy (anonymous):

2 sides are equal

OpenStudy (anonymous):

AK= 3x + 1 KB = 1/2x+6 ak = kb correct?

OpenStudy (anonymous):

its like 3 in the morning where i am so im not relally into it all so as much info given is the better

OpenStudy (anonymous):

y are you up 3 in the morning?

OpenStudy (anonymous):

because i have to get my school done

OpenStudy (anonymous):

ne way AK = KB

OpenStudy (anonymous):

i have like 9 more questions that i need help on

OpenStudy (anonymous):

AK= 3x + 1 KB = 1/2x+6 3x + 1 = = 1/2x+6 ^ ^ AK KB this is the equation you have to solve

OpenStudy (anonymous):

you see how it works

OpenStudy (anonymous):

but that doesnt give me AK

OpenStudy (anonymous):

ya it does

OpenStudy (anonymous):

because thats what i did to get AB in the previous question

OpenStudy (anonymous):

if you get AB and you want AK K is the midpoint ( bisector = splits line in half)

OpenStudy (anonymous):

and its 1/3x+9 for this

OpenStudy (anonymous):

this is just an equation

OpenStudy (anonymous):

okies lemme work this out

OpenStudy (anonymous):

Find the slope of a line that has an x-intercept of -2 and a y-intercept of 6

OpenStudy (anonymous):

State whether the slope of a line passing through (2, -3) and (1, 8) is positive, negative, zero or undefined

OpenStudy (anonymous):

Find the value of y so that the line passing through (6, y) and (7, 5) has a slope of -3

OpenStudy (anonymous):

Determine if the lines that pass through the given points are parallel, perpendicular or neither. Line A: (5, 10) and (12, 6) Line B: (-2, 4) and (5, 8)

OpenStudy (anonymous):

Determine if the lines that pass through the given points are parallel, perpendicular or neither. Line A: (2, 3) and (--4, 5) Line B: (-4, 9) and (-3, 12)

OpenStudy (anonymous):

Find the rate of change between the two points: (11, 250) and (16, 200) where x is in days and y is in dollars

OpenStudy (anonymous):

Write the equation of a line that passes through the points ( -2, 5) and (-2, 11)

OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that has a slope of -1/4 and passes through the point (8, -1

OpenStudy (anonymous):

Write the equation of a line in slope-intercept form passing through (2, 7) and (-2, 9

OpenStudy (anonymous):

can you help me with these?

OpenStudy (anonymous):

trying

OpenStudy (anonymous):

thank you <3

OpenStudy (anonymous):

do you know how to plot points on a graph?

OpenStudy (anonymous):

not realy

OpenStudy (anonymous):